OpenStudy (anonymous):

This sequence is geometric. What is the next term? 500, 250, 125, 62.5, ... Answer 31.25 12.5 25 250

OpenStudy (anonymous):

You might notice that since it is geometric, the ratios are same, for example,\[\frac{ 500 }{ 125}=\frac{ 250 }{ 125 }=\frac{ 125 }{ 62.5 }=\frac{ 62.5 }{ x }\] I believe you can get x.

OpenStudy (anonymous):

The answer is 12.5 correct @Shadowys

OpenStudy (anonymous):

i made a mistake there. it's 500/250 = 250/125. sorry about that.

OpenStudy (anonymous):

So 250 than? thanks

OpenStudy (anonymous):

u got it right why did u delete it?

OpenStudy (anonymous):

yh. You should be able to get it. lol

OpenStudy (anonymous):

So 250 is the answer thanks

OpenStudy (anonymous):

no 62.5/2=?

OpenStudy (anonymous):

31.25

OpenStudy (anonymous):

yup!